Grade 12th passAlgebra

The measures of two adjacent angles of a parallelogram are in the ratio 3: 2. Find the measure of each of the angles of the parallelogram.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let the measures of two adjacent angles ∠A and ∠B be 3x and 2x respectively inparallelogram ABCD.
∠A+∠B=180°
⇒3x+2x=180°
⇒5x=180°
⇒x=36°
We know that opposite sides of a parallelogram are equal.
∠A = ∠C = 3x = 3 × 36° = 108°
∠B = ∠D = 2x = 2 × 36° =72°

Thanks